# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v1

@@ VI.2 Reversible asymptotic conversion: Proof of the Marvian-Spekkens conjecture
[…]
  1. (i)
    $\mathrm{Sym}_{G}(\psi)=\mathrm{Sym}_{G}(\phi)$.
  2. (ii)
    $C_{\mathfrak{g}}(\psi)=R(\psi\to\phi)C_{\mathfrak{g}}(\phi)$.
  3. (iii)
[…]
VI.3 Irreversibility in asymptotic conversion
$\mathcal{M}(\psi)$ $\coloneqq 2\braket{\psi}{\mathcal{J}}{\psi},$ (47)
[…]
Since the QGT is calculated for $\psi$ as
VI.4 Finite groups
reproducing the result on the approximate asymptotic conversion rate in [34].
The proof of the direct part provides a clear understanding of the reason why the approximate asymptotic conversion rate diverges for finite groups. Since there is only a finite number of different elements in $\{\mathcal{U}_{g}(\psi)\mid g\in G\}$, it is possible to identify the state $\mathcal{U}_{g}(\psi)$ with an exponentially small failure probability using a state tomography protocol on $\mathcal{U}_{g}(\psi)^{\otimes N}$. Let $\hat{g}\in G$ denote the estimate of the true value $g$ obtained by this tomography protocol. Depending on the estimate $\hat{g}$, we can prepare an arbitrary number of copies of $\mathcal{U}^{\prime}_{\hat{g}}(\phi)$. The condition $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ implies that if $\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ is satisfied, then $\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}^{\prime}_{g}(\phi)$ also holds. Therefore, the approximate asymptotic conversion rate diverges since it is possible to create an arbitrary number of copies of $\mathcal{U}^{\prime}_{g}(\phi)$ from $\mathcal{U}_{g}(\psi)^{\otimes N}$ with asymptotically vanishing failure probability. We emphasize that the exact asymptotic conversion rate in [34] is not covered by this analysis because there always remains an error in the tomography process.
VII Conclusion
@@
In this paper, we established the approximate asymptotic conversion theory for i.i.d. pure states in RTA for a compact Lie group. Our result provides a formula for the asymptotic conversion rate in RTA for any compact Lie group by using QGTs, implying that QGT fulfills the same function in RTA as entanglement entropy does in entanglement theory. In other words, the present study unveils the distinguished role of QGT as an asymmetry measure governing the conversion rate in resource manipulation in RTA. As a demonstration, we applied our formula to several different setups. Prior to the present research, the i.i.d. asymptotic conversion theory among pure states in RTA for a Lie group is established only for $U(1)$ group [21], and a few examples of restricted sets of pure states have been studied for other Lie groups [2, 37]. We first proved that the formula for the conversion rate in RTA for $U(1)$ in [21] is reproduced from our formula. We also affirmatively proved the Marvian-Spekkens conjecture [6, 72] on reversible conversion in RTA as a corollary of our main theorem, which has remained unproven for a decade. We also studied the examples where the asymptotic conversion is irreversible for $U(1)$ and $SU(2)$ groups. The conversion rate calculated for this example of $SU(2)$ is consistent with the result in [2], where the conversion rate is studied on a restricted set of pure states. Furthermore, we demonstrated that RTA for finite groups can also be analyzed in our framework by appending a trivial phase, providing an alternative proof of the divergence in the approximate asymptotic conversion rate for any finite group [34].
@@ Acknowledgements.
The authors would like to thank Tomohiro Shitara for providing his expert knowledge on monotone metrics, Yui Kuramochi for valuable comments on an early draft of this manuscript, and also Iman Marvian for the valuable discussion. K.Y. acknowledges support from JSPS KAKENHI Grant No. JP24KJ0085. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. H.T. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061.
@@ Appendix A Proof for the main theorem for any projective unitary representations
In the main text, we proved Eq. (7) for (non-projective) unitary representations of a compact Lie group. In Appendix A.1, we first review the result in [34] that relates the conversion rates for projective unitary representations to that for (non-projective) unitary representations and explain a formula for the conversion rate for projective unitary representations that are continuous but not differentiable. In Appendix A.2, we first show that the QGTs at different points of a connected component of a Lie group are interrelated by a congruence transformation, which plays an essential role in proving Theorem 1 from Eq. (7). In Appendix A.3, we complete the proof for Eq. (5) in Theorem 1 and Eq. (7) for any projective unitary representations that are differentiable. Furthermore, Eq. (8) is proven Sec. A.4.
@@ Appendix B Proofs of lemmas in the converse part
B.1 Proof of the equivalence between two alternative definitions of asymptotic convertibility

### v2

@@ VII.1 Distillation of asymmetry
Suppose $\{\rho^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor rN\rfloor}\}_{N}$ for $r\geq 0$. From Eq. (32) and Lemma 4, as an extension of Eq. (34), we get
$f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}\geq\frac{M}{N}\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-\frac{M}{N}h(\epsilon)+\frac{1}{N}o(M),$ (71)
where $M\coloneqq\lfloor rN\rfloor$. In the limit of $N\to\infty$, we have
$f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}^{2}\geq r\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-rh(\epsilon).$ (72)
Since this inequality holds for all sufficiently small $\epsilon>0$, we get $f_{q}(0)\|\mathrm{i}[\bm{\gamma}^{\dagger}\bm{X},\rho]\|_{f_{q},\rho}\geq r\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}$. In the limit of $q\to 1^{-}$, we obtain $\mathcal{S}^{\rho}\geq r\mathcal{Q}^{\phi}$. Repeating this argument for $\mathcal{U}_{g}(\rho)$ and $\mathcal{U}^{\prime}_{g}(\phi)$ instead of $\rho$ and $\phi$, we get
$\forall g\in G,\quad\mathcal{S}^{\mathcal{U}_{g}(\rho)}\geq r\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)},$ (73)
which concludes the proof of Eq. (70). ∎
This result first establishes an upper bound on the distillable asymmetry, which applies to any state $\rho$ and any compact Lie group $G$, and becomes tight when $\rho$ is pure.
We remark that in a special case where $[\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ holds for some $\bm{\gamma}\in\mathbb{C}^{\dim G}$, we have $\bm{\gamma}^{\dagger}\mathcal{S}^{\rho}\bm{\gamma}=0$. As an immediate corollary of the above theorem, we therefore get the following:
Corollary 13 (Sufficient condition for vanishing distillable asymmetry).
If there exists a vector $\bm{\gamma}\in\mathbb{C}^{\dim G}$ satisfying $[\Pi_{\rho},\bm{\gamma}^{\dagger}\bm{X}]=0$ and $\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}\neq 0$, then $\mathcal{A}_{\mathrm{d}}(\rho)=0$.
This corollary implies that for a typical state, which has full rank, $\mathcal{A}_{\mathrm{d}}(\rho)=0$ holds for any asymmetric pure state $\phi$. This result is consistent with a result in prior research [19] on time-translation asymmetry, where the distillable coherence, defined as the optimal conversion rate from a state $\rho$ to a pure state $\phi$ given by a superposition of the ground and excited state of a Hamiltonian $H$, is shown to be zero when $[\Pi_{\rho},H]=0$. We remark that although a state $\rho$ satisfying the condition in Corollary 13 has zero distillable asymmetry, it does not necessarily implies that $\rho$ is useless in asymptotic conversion. For example, while a full-rank state has zero distillable asymmetry, a sublinear number of such states may be used to enhance the conversion rate by removing the restriction imposed by symmetry subgroups in asymptotic conversion among pure states, as stated in Theorem 11.
VII.2 Dilution of asymmetry
Next, we investigate the opposite scenario, where we create a general state $\rho$ by consuming the reference pure state $\phi$. The asymmetry cost is defined as
$\mathcal{A}_{\mathrm{c}}(\rho)\coloneqq\inf\{r\geq 0\mid\{\phi^{\otimes\lceil rN\rceil}\}_{N}\xrightarrow{G\text{-cov.}}\{\rho^{\otimes N}\}_{N}\},$ (74)
i.e., $\mathcal{A}_{\mathrm{c}}(\rho)=1/R(\rho\to\phi)$, where $1/\infty$ and $1/0$ are formally regarded as 00 and $\infty$, respectively. We remark that Eq. (8) implies $\mathcal{A}_{\mathrm{c}}(\rho)<\infty$ only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$. Thus, we only investigate $\mathcal{A}_{\mathrm{c}}(\rho)$ for $\rho$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho)$.
Following the standard argument using typical sequence [72, 21], it is proven that
$\sum_{i=1}^{k}p_{i}\mathcal{A}_{\mathrm{c}}(\rho_{i})\geq\mathcal{A}_{\mathrm{c}}(\rho),$ (75)
where $\{\rho_{i}\}_{i=1}^{k}$ is a set of states and $\{p_{i}\}_{i=1}^{k}$ is a probability distribution such that $p_{i}>0$ and $\sum_{i=1}^{k}p_{i}=1$. An intuition behind this inequality is that, instead of preparing $\rho$ directly, we can adopt a strategy where $\{\rho_{i}\}_{i=1}^{k}$ is probabilistically generated according to the probability distribution $\{p_{i}\}_{i=1}^{k}$, although this strategy is not necessarily optimal. See Appendix E for the proof of Eq. (75). We remark that the left-hand side of Eq. (75) is finite only if $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\rho_{i})$ for all $i=1,\cdots,k$.
In light of the pure-state conversion theory established in Theorem 1, a promising approach to obtaining a computable upper bound of $\mathcal{A}_{\mathrm{c}}(\rho)$ is to decompose a state $\rho$ into an ensemble of pure states. Suppose that a state $\rho$ can be decomposed as $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$, where a symmetric (possibly mixed  [^2] We remark that not all symmetric mixed states can be decomposed as an ensemble of symmetric pure states. Indeed, when a unitary representation of a group is decomposed into irreducible components, Schur’s lemma implies that only a one-dimensional irreducible representation contains a symmetric pure state. In other words, no symmetric pure state exists in any higher-dimensional irreducible representation.) state $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i=1}^{k-1}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$. Then, Eq. (7) implies
$\sum_{i=1}^{k-1}p_{i}r_{i}\geq\mathcal{A}_{\mathrm{c}}(\rho),$ (76)
where
$r_{i}\coloneqq\inf\left\{r\geq 0\,\middle|\,\forall g\in G,\,r\mathcal{Q}^{\mathcal{U}_{g}(\phi)}\geq\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\psi_{i})}\right\}.$ (77)
When a state $\rho$ can be obtained as an ensemble of pure states and symmetric state, minimizing the left-hand side of Eq. (76) over such an ensemble leads to a tighter bound on the asymmetry cost $\mathcal{A}_{\mathrm{c}}(\rho)$. In the case of $G=U(1)$, it is known that such a minimization yields a tight bound [21]. Let $U(\theta)=e^{\mathrm{i}H\theta}$ and $U^{\prime}(\theta)=e^{\mathrm{i}H^{\prime}\theta}$ be unitary representations of $G=U(1)$ on finite-dimensional Hilbert spaces with Hamiltonians $H=\sum_{n=0}^{d-1}n\ket{n}\bra{n}$ | and $H^{\prime}=\sum_{n=0}^{d^{\prime}-1}n\ket{n}\bra{n}$ |. We fix the reference state $\phi$ to be the so-called coherence bit, given by |$\ket{\phi}=(\ket{0}+\ket{1})/\sqrt{2}$ and assume that $\mathrm{Sym}_{G}(\rho)=\mathrm{Sym}_{G}(\phi)$. Since $r_{i}=V(\psi_{i},H^{\prime})/V(\phi,H)=4V(\psi_{i},H^{\prime})$, Eq. (76) implies $\min_{\{p_{i},\psi_{i}\}_{i}}4\sum_{i}p_{i}V(\psi_{i},H^{\prime})\geq\mathcal{A}_{\mathrm{c}}(\rho)$, where the minimization is taken over the set of all pure-state ensembles $\{p_{i},\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\psi_{i})$ and $\rho=\sum_{i}p_{i}\psi_{i}$. Note that it suffices to consider pure-state ensembles for $G=U(1)$ since any symmetric state is given as a probabilistic mixture of symmetric pure states. It was conjectured in [74] and later proven in [75] that $\min_{\{p_{i},\psi_{i}\}_{i}}4\sum_{i}p_{i}V(\psi_{i},H^{\prime})=\mathcal{F}_{H^{\prime}}(\rho)$, where $\mathcal{F}_{H^{\prime}}(\rho)$ denotes the SLD quantum Fisher information, given by $\mathcal{F}_{H^{\prime}}(\rho)\coloneqq 2\sum_{i,j}\frac{(\lambda_{i}-\lambda_{j})^{2}}{\lambda_{i}+\lambda_{j}}|\braket{i}{H}{j}|^{2}$ using the eigenvalue decomposition $\rho=\sum_{i}\lambda_{i}\ket{i}\bra{i}$ |. Furthermore, in [21], it was shown that the minimum is attained by a set of pure states $\{\psi_{i}\}_{i}$ such that $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\psi_{i})$, which implies $\mathcal{F}_{H^{\prime}}(\rho)\geq\mathcal{A}_{\mathrm{c}}(\rho)$. In [21], the opposite inequality was also shown by proving the monotonicity of the SLD quantum Fisher information rate under asymptotic conversion, which completed the proof of $\mathcal{F}_{H^{\prime}}(\rho)=\mathcal{A}_{\mathrm{c}}(\rho)$ for $G=U(1)$. In contrast to the $U(1)$ case, however, we remark that decomposition into $\rho=\sum_{i=1}^{k-1}p_{i}\psi_{i}+p_{\mathrm{s}}\rho_{\mathrm{s}}$ with symmetric $\rho_{\mathrm{s}}$ and pure states $\{\psi_{i}\}_{i}$ satisfying $\mathrm{Sym}_{G}(\phi)\subset\mathrm{Sym}_{G}(\psi_{i})$ does not always exist for a general Lie group $G$. See Appendix E for a concrete example for $G=U(4)$.
Given the success in the case of $G=U(1)$, where the property of the SLD quantum Fisher information $\mathcal{F}_{H}(\rho)=\min_{\{p_{i},\psi_{i}\}_{i}}4\sum_{i}p_{i}V(\psi_{i},H)$ is utilized, a natural extension is to consider the convex roof of QGT, namely $\min_{\{p_{i},\psi_{i}\}_{i}}\sum_{i}p_{i}\mathcal{Q}^{\psi_{i}}$, as a possible asymmetry quantifier in RTA for a general compact Lie group. However, as we demonstrate in Appendix F, $\min_{\{p_{i},\psi_{i}\}_{i}}\sum_{i}p_{i}\mathcal{Q}^{\psi_{i}}$ does not always exist. Analogous to another property of the SLD quantum Fisher information, $\mathcal{F}_{H}(\rho)=\min_{\Psi_{\rho},H_{A}}4V(\Psi_{\rho},H+H_{A})$, where the minimization is taken over any purification $\Psi_{\rho}$ of $\rho$ and Hermitian operators $H_{A}$ on the ancillary system [21], we also examine the minimization of the QGT of purifications of a state $\rho$. We show that such a minimum does not also exist in general in Appendix G. The non-existence of these extensions of the SLD quantum Fisher information stems from the fact that a partial order on matrices is not necessarily a total order, highlighting the difference between Abelian and non-Abelian symmetries.
VIII Conclusion
@@
In this paper, we established the approximate asymptotic conversion theory for i.i.d. pure states in RTA for a compact Lie group. Our result provides a formula for the asymptotic conversion rate in RTA for any compact Lie group by using QGTs, implying that QGT fulfills the same function in RTA as entanglement entropy does in entanglement theory. In other words, the present study unveils the distinguished role of QGT as an asymmetry measure governing the conversion rate in resource manipulation in RTA. As a demonstration, we applied our formula to several different setups. Prior to the present research, the i.i.d. asymptotic conversion theory among pure states in RTA for a Lie group is established only for $U(1)$ group [21], and a few examples of restricted sets of pure states have been studied for other Lie groups [2, 37]. First, we proved that our formula reproduces the formula for the conversion rate in RTA for $U(1)$ in [21]. Next, we affirmatively proved the Marvian-Spekkens conjecture [6, 76] on reversible conversion in RTA as a corollary of our main theorem, which has remained unproven for a decade. We also studied the examples where the asymptotic conversion is irreversible for $U(1)$ and $SU(2)$ groups. The conversion rate calculated for this example of $SU(2)$ is consistent with the result in [2], where the conversion rate is studied on a restricted set of pure states. Furthermore, we demonstrated that RTA for finite groups can also be analyzed in our framework by appending a trivial phase, providing an alternative proof of the divergence in the approximate asymptotic conversion rate for any finite group [34]. Additionally, we examined an alternative setup in which a sublinear number of states is available as an auxiliary resource, which removes the constraints originating from symmetry subgroups and thereby improves the conversion rate.
While this paper established an i.i.d. pure-state conversion theory in the RTA for a compact Lie group, a comprehensive understanding of mixed-state asymmetry remains elusive. We investigated both the distillation and dilution of asymmetry, i.e., the conversion of a mixed state to and from a pure reference state. In the former scenario, we derived an upper bound on distillable asymmetry by introducing a new asymmetry measure, which can be viewed as an extension of QGT to mixed states. In the latter scenario, we pointed out that a strategy for creating a mixed state as a pure-state ensemble, which is shown to be optimal in the $U(1)$ case, is not always successful for other Lie groups. Although these findings are not exhaustive, we speculate that they lay a foundation for future investigations.
@@ Acknowledgements.
The authors would like to thank Yui Kuramochi for valuable comments on an early draft of this manuscript, and also Iman Marvian for the valuable discussion. K.Y. acknowledges support from JSPS KAKENHI Grant No. JP24KJ0085. Y.M. is supported by JSPS KAKENHI Grant No. JP23KJ0421. T.S. acknowledges support from JST Moonshot R&D Grant No. JPMJMS2061, JST CREST Grant No. JPMJCR23I4, and MEXT Q-LEAP Grant No. JPMXS0120319794, Japan. H.T. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061.
@@ Appendix A Proof for the main theorem for any projective unitary representations
In the main text, we proved Eq. (7) for (non-projective) unitary representations of a compact Lie group. In Appendix A.1, we first review the result in [34] that relates the conversion rates for projective unitary representations to that for (non-projective) unitary representations and explain a formula for the conversion rate for projective unitary representations that are continuous but not differentiable. In Appendix A.2, we show that the QGTs at different points of a connected component of a Lie group are interrelated by a congruence transformation, which plays an essential role in proving Theorem 1 from Eq. (7). In Appendix A.3, we complete the proof for Eq. (5) in Theorem 1 and Eq. (7) for any projective unitary representations that are differentiable. Furthermore, Eq. (8) is proven Sec. A.4.
@@ Appendix B Proofs of lemmas in the converse part
B.1 Proof of Lemma 3
In the literature, the monotonicity of a monotone metric is often proven under the assumption that $\rho$ and $\mathcal{E}(\rho)$ are invertible. We here explicitly prove the monotonicity of a monotone metric for an operator monotone function $f_{q}(x)=(1-q)+qx$ for $q\in(0,1)$ without assuming states are invertible for completeness. For this purpose, we here introduce several notations in [68]. Let $\mathcal{M}(\mathcal{H})$ be the set of all linear operators on a Hilbert space $\mathcal{H}$. For an arbitrary probability distribution $p$ on $[0,1]$, we define a linear map $E_{p,\rho}:\mathcal{M}(\mathcal{H})\to\mathcal{M}(\mathcal{H})$ by
$E_{p,\rho}(A)$ $\coloneqq\int_{0}^{1}\rho^{\lambda}A\rho^{1-\lambda}p(\mathrm{d}\lambda).$ (107)
Note that when $\rho$ is not invertible, the map $E_{p,\rho}$ has a non-trivial kernel. We denote the image of $E_{p,\rho}$ by $\mathcal{M}_{p,\rho}^{(m)}(\mathcal{H})$. For operators $A,B\in\mathcal{M}_{p,\rho}^{(m)}(\mathcal{H})$, we define $\braket{A,B}_{p,\rho}^{(m)}=\mathrm{Tr}\left(C^{\dagger}B\right)$, where $C\in\mathcal{M}(\mathcal{H})$ is an arbitrary operator satisfying $E_{p,\rho}(C)=A$. This inner product is well defined since for operators $C,C^{\prime}\in\mathcal{M}(\mathcal{H})$ satisfying $E_{p,\rho}(C)=E_{p,\rho}(C^{\prime})=A$, we get $\mathrm{Tr}(C^{\dagger}B)=\mathrm{Tr}(C^{\dagger}E_{p,\rho}(D))=\mathrm{Tr}((E_{p,\rho}(C))^{\dagger}D)=\mathrm{Tr}((E_{p,\rho}(C^{\prime}))^{\dagger}D)=\mathrm{Tr}(C^{\prime{\dagger}}B)$, where $D\in\mathcal{M}(\mathcal{H})$ is an operator satisfying $E_{p,\rho}(D)=B$.
As shown in Theorem 6.1 in [68], the induced norm $\|A\|^{(m)}_{p,\rho}\coloneqq\sqrt{\braket{A,A}_{p,\rho}^{(m)}}$ is monotonically non-increasing under any CPTP map $\kappa$. That is,
$\|A\|_{p,\rho}^{(m)}\geq\|\kappa(A)\|_{p,\kappa(\rho)}^{(m)}$ (108)
holds for any operator $A$ such that $A\in\mathcal{M}_{p,\rho}^{(m)}(\mathcal{H})$ and $\kappa(A)\in\mathcal{M}_{p,\kappa(\rho)}^{(m)}(\mathcal{H})$.
In order to relate the above notation with ours, let us now consider a particular probability distribution $p_{q}\coloneqq(1-q)\delta_{1}+q\delta_{0}$ characterized by $q\in(0,1)$, where $\delta_{0}$ and $\delta_{1}$ denote the Dirac measure concentrated on 00 and $1$, respectively. In this case, we have
$E_{p_{q},\rho}(A)=(1-q)\rho A+qA\rho=m_{f_{q}}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(A),$ (109)
where $m_{f}(x,y)\coloneqq yf(x/y)$ and $f_{q}$ is an operator monotone function defined in Eq. (20). Then, we have the following:
Lemma 19.
For any state $\rho$ and linear operator $B$, $\mathrm{i}[\rho,B]\in\mathcal{M}_{p_{q},\rho}^{(m)}(\mathcal{H})$.
Proof.
Let $\rho=\sum_{i=1}^{d}p_{i}\ket{i}\bra{i}$ | be the eigenvalue decomposition of $\rho$. For an operator $A$ defined by
$A\coloneqq\sum_{k,l;m_{f_{q}}(p_{k},p_{l})>0}\frac{1}{m_{f_{q}}(p_{k},p_{l})}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k},$ (110)
we get
$E_{p_{q},\rho}(A)$ $=\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\frac{(1-q)p_{l}}{(1-q)p_{l}+qp_{k}}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}+\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\frac{qp_{k}}{(1-q)p_{l}+qp_{k}}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (111)
$=\sum_{k,l;(1-q)p_{l}+qp_{k}>0}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (112)
$=\mathrm{i}[\rho,B]-\sum_{k,l;p_{l}=p_{k}=0}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}$ | (113)
$=\mathrm{i}[\rho,B],$ (114)
where we have used the fact that $(1-q)p_{l}+qp_{k}>0$ unless $p_{k}=p_{l}=0$. ∎
Therefore, for any state $\rho$ and linear operator $B$, we get
$\left(\|\mathrm{i}[\rho,B]\|_{p_{q},\rho}^{(m)}\right)^{2}$ $=\mathrm{Tr}\left(\left(\sum_{k,l;m_{f_{q}}(p_{k},p_{l})>0}\frac{1}{m_{f_{q}}(p_{k},p_{l})}\ket{l}\braket{l}{\mathrm{i}[\rho,B]}{k}\bra{k}\right)^{\dagger}(\mathrm{i}[\rho,B])\right)$ (115)
$=\braket{\mathrm{i}[\rho,B],\mathrm{i}[\rho,B]}_{f_{q},\rho},$ (116)
where we used Eq. (18) in the last equality. Using this relation, the monotonicity in Eq. (108) completes the proof of Lemma 3.
B.2 Proof of the equivalence between two alternative definitions of asymptotic convertibility